# Hyperflux: Pruning Reveals the Importance of Weights

### v2

@@ Abstract
Network pruning is used to reduce inference latency and power consumption in large neural networks. However, most existing methods use ad-hoc heuristics, lacking much insight and justified mainly by empirical results. We introduce Hyperflux, a conceptually-grounded $L_{0}$ pruning approach that estimates each weight’s importance through its flux, the gradient’s response to the weight’s removal. A global pressure term continuously drives all weights toward pruning, with those critical for accuracy being automatically regrown based on their flux. We postulate several properties that naturally follow from our framework and experimentally validate each of them. One such property is the relationship between final sparsity and pressure, for which we derive a generalized scaling-law equation that is used to design our sparsity-controlling scheduler. Empirically, we demonstrate state-of-the-art results with ResNet-50 and VGG-19 on CIFAR-10 and CIFAR-100.
@@ 1 Introduction
Overparameterization has become the norm in modern deep learning to achieve state-of-the-art performance Neyshabur et al. (2019); Allen-Zhu et al. (2019); Li et al. (2018). Despite clear benefits for training, this practice also increases computational and memory costs, complicating deployment on resource-constrained devices such as edge hardware, IoT platforms, and autonomous robots (Shi et al., 2016; Li et al., 2019). Recent theoretical and empirical findings suggest that sparse subnetworks extracted from large dense models can match or exceed the accuracy of their dense counterparts Frankle and Carbin (2019); Zhou et al. (2019); Ma et al. (2021); Lee et al. (2019); De Jorge et al. (2021); Cho et al. (2023); Yite et al. (2023); Frantar et al. (2024); Wang et al. (2023) and even outperform smaller dense models of equal size Ramanujan et al. (2020); Li et al. (2020); Zhu and Gupta (2018). These results have created interest in network pruning as a strategy to identify minimal, high-performing subnetworks.
@@
Pruning has a rich history LeCun et al. (1989); Mozer and Smolensky (1988); Thimm and Hoppe (1995) and continues to prove valuable for real-time applications Han et al. (2016); Jongsoo et al. (2017); Wang et al. (2019). Recent methods have significantly advanced the field by resorting to a variety of strategies and heuristics, from magnitude pruning, gradient methods, and Hessian-based criteria Han et al. (2015, 2016); LeCun et al. (1992); Singh and Alistarh (2020); Bellec et al. (2018); Frankle and Carbin (2019) to dynamic pruning approaches Liu et al. (2020); Cho et al. (2023); Savarese et al. (2020); Kusupati et al. (2020); Wortsman et al. (2019) or combinations thereof Liu et al. (2022); Evci et al. (2020). However, the strong interdependence between weights remains a challenge Jin et al. (2020); Templeton et al. (2024); Lee et al. (2019); De Jorge et al. (2021); Louizos et al. (2017), as it complicates the task of determining each weight’s importance. Optimal pruning has been explored Hoefler et al. (2021); LeCun et al. (1992), but such formulations are typically computationally intractable in practice. In contrast, most current state-of-the-art strategies prioritize empirical results and speed through heuristics, at the expense of theoretical grounding.
Given this gap, we ask: Can we create a pruning method that is both empirically strong and conceptually grounded?
@@
Inspired by the principle that the value of something is not truly known until it is lost, which has shaped major discoveries in fields such as functional genomics Giaever et al. (2002); Shalem et al. (2014), neuroscience Rorden and Karnath (2004), and network science Albert et al. (2000), we introduce Hyperflux, an $L_{0}$ pruning method that determines a weight’s importance by first removing it. Unlike most works, Hyperflux puts a large emphasis on conceptual grounding and explainability.
@@
The main idea of our method is that each weight has a flux, which appears when the weight is pruned through the network’s gradients. A global $L_{0}$ regularization term called pressure pushes all weights towards pruning, aiming to uncover each of their fluxes. Those weights whose flux is greater than the pressure will be regrown, while the rest will remain pruned. This process is repeated until the end of training. A useful side effect of pruning and regrowth happening concurrently on all weights multiple times is that the network’s topology implicitly becomes noisy, disentangling the overall weight evaluation from a specific topology.
We postulate several properties that emerge from our framework: sparsity convergence, a sparsity-pressure relationship, and large flux for important weights. We empirically confirm each of these properties and, for the sparsity-pressure relationship, we obtain dependencies similar to those of known scaling laws in neural networks Hestness et al. (2017); Kaplan et al. (2020); Tom et al. (2020); Rosenfeld et al. (2020); Gordon et al. (2021); Hernandez et al. (2021); Zhai et al. (2021); Hoffmann et al. (2022). Based on the postulated properties, we propose a pressure scheduler, as well as a stabilization stage after pruning, further differentiating Hyperflux from recent $L_{0}$ methods Louizos et al. (2018); Savarese et al. (2020); Yamada et al. (2020); Liu et al. (2020). The scheduler is used to achieve the desired sparsity, after which the stabilization stage recovers accuracy lost to noise induced by pruning.
@@
  • We introduce Hyperflux, a conceptually grounded pruning method which develops the notions of flux and pressure, before empirically studying their emergent properties.
  • Based on these properties, we introduce a pressure-controlling scheduler to achieve a desired sparsity, as well as a stabilization stage after pruning.
  • We obtain state-of-the-art results, achieving better or comparable accuracy to existing methods in empirical validation across several networks and datasets.
@@ 3.2 Weight flux
Figure 2: Depiction of gradients (as arrows) influencing $t_{i}$, red, yellow and blue denote pressure, $\mathcal{G}_{i}^{+}$ and respectively $\mathcal{G}_{i}^{-}$.
$\sum_{h=P_{s}}^{P_{f}}(\mathcal{G}_{i}(\omega,\mathcal{T}^{h})-\frac{\gamma}{d})>0\iff(P_{f}-P_{s})\cdot\Bigl{[}\mathcal{G}_{i}^{-}\!\Bigl{(}\omega,\mathcal{T}^{P_{s}\to P_{f}}\Bigr{)}-\frac{\gamma}{d}\Bigr{]}>0.$ (6)
3.3 Pressure & Flux Properties
@@
Following from the theoretical insights about flux and pressure described so far, we postulate a series of properties that naturally emerge from these concepts. We experimentally validate each one of the properties, confirming our expectations, and laying the foundation for our $\gamma$ scheduler.
@@
Property 1: Sparsity Convergence for a Fixed $\gamma$. As sparsity increases and the number of weights decreases, fewer weights are used to represent the same information contained within the dataset, so the overall importance and flux of the remaining weights should be larger. Once the flux of the remaining weights surpasses the pressure, sparsity should converge. Therefore, we ask the following question: Given a fixed $\gamma$, will the network converge to a final sparsity $\mathcal{S}$? In Figure 3(a), we test this by running LeNet-300 on MNIST and ResNet-50 on Cifar-10. We allow each network to train for 300 to 1000 epochs with a constant pressure $\gamma$ and observe the results.
Figure 3: Convergence for fixed $\gamma=2$ is showcased in (a), while in (b) we present the relationship between $\gamma$ and final sparsity.
We test two different optimizers for $t$ values, SGD and Adam, while for weights we use the same Adam optimizer everywhere (more on training setup and its notation in Appendix F). Our findings suggest that there is no one curve that fits the decrease in parameters for both optimizers, but $\mathcal{S}$ is the same regardless of the optimizer used. An important observation is that $\mathcal{S}$ is influenced by the weights learning rate $\eta_{\omega}$. If $\eta_{\omega}$ is high, convergence happens in a larger number of epochs (1000 in our experiments), at a higher sparsity. If $\eta_{\omega}$ is low, convergence happens sooner (300 epochs), at a lower sparsity. One way to ensure smooth convergence is to decrease $\eta_{\omega}$ during training. Otherwise, the network tends to converge more slowly, as seen in the green curve experiment. Further ablation studies are found in Appendix B.
Property 2: Relationship Between $\gamma$ and Final Sparsity. Assuming as illustrated above that all networks have a sparsity they converge to for a fixed $\gamma$, we ask: Can we find the relationship between $\gamma$ and $\mathcal{S}$? We modify the previous experiment to run the networks for 300 epochs with the same training setup for several values of $\gamma$. Our empirical results suggest a generalized scaling law:
$\ln(s)\;=\;\ln(c)\;-\;\alpha_{0}\,\ln(\gamma)\;-\;\alpha_{1}\,\bigl{(}\ln(\gamma)\bigr{)}^{2}$ (7)
where constants, c, $\alpha_{0}$, $\alpha_{1}$ depend on dataset, network architecture and training setup. Figure 3(b) showcases different convergence points for different optimizers and $\gamma$ values. The curves bend more sharply toward the end as the network loses accuracy (and feature representations), yielding a lower convergence point, until the network collapses, pruning all weights. We call this property by the name Neural Pruning Law.
Table 1: Pressure needed to prune the bottleneck
Sparsity (%) Pressure
99.05 $\gamma_{B}$
99.75 $\gamma_{B}\cdot 2^{3}$
99.95 $\gamma_{B}\cdot 2^{7}$
Bottleneck $\gamma_{B}\cdot 2^{13}$
Property 3: Important weights developing large flux is probably the most important idea in Hyperflux. Therefore we ask: How large is the flux of critical weights compared to other weights? To obtain a set of critical weights, we create a bottleneck in a LeNet-300 network by pruning only the last weight matrix, until an identity remains between the hidden layer and each unit of the output layer (in our case 10 weights). We measure their flux by reporting the largest pressure that still does not prune the weights, since we know that weights carrying greater flux demand higher pressure to prune, see equation (6).
As a baseline of comparison, we use the original network without the bottleneck, and report several $\gamma$ and the corresponding $\mathcal{S}$ they produce. The results are reported in Table 1. The bottleneck weights have $2^{13}$ more flux than the weights of a $99\%$ sparsified network.
@@ 3.4 Pressure Scheduler & Stabilization Stage
Our findings from Section 3.3 suggest that a $\gamma$ exists for any desired $\mathcal{S}$. However, in practical applications, $\gamma$ is not known at the start and tuning it would require hyperparameter search. Instead, we propose a dynamic scheduler that adjusts $\gamma$ after each epoch automatically, driving the network towards a desired sparsity. Furthermore, to ensure convergence of weights after pruning, we introduce a stabilization stage at the end.
@@
1:Input: Current sparsity $s_{e}$ and epoch $e$
2:Requires: Pruning epochs $E_{p}$, desired final sparsity $\mathcal{S}$, pressure policy $\Pi$, step $u$, exponent $\alpha$.
@@
Pressure Scheduler: The goal of our scheduler is to adjust $\gamma$ such that the network converges to $\mathcal{S}$ with minimal accuracy decrease. We denote by $\gamma_{e}$ and $s_{e}$, the pressure and network sparsity at epoch $e$. Because the frequency of updates is constant, occurring after each epoch, any non-linear change in pressure required to affect sparsity (see Eq. 7) must arise from the update rule. To get this nonlinearity we set $\gamma_{e}=(p_{e})^{\alpha},$ with $p_{e}$, a scalar base, updated according to Algorithm 1. Inertia terms $p_{+}$ and $p_{-}$ account for suboptimal $\alpha$ or $u$. Apart from non-linear updates, our scheduler requires a binary pressure policy $\Pi$ to determine when those updates are applied such that $\mathcal{S}$ is reached. We explore two choices for $\Pi$. In the first, $s_{e}$ follows a user-defined curve $f(e)$, trading precise convergence to $\mathcal{S}$ for trajectory control. In this case $\Pi(E_{p},\mathcal{S},s_{e},e)$ is true if and only if $s_{e}<f(e)$. In the second policy, $s_{e}$ stays between a dynamic upper bound and the target $\mathcal{S}$, achieving precise convergence to $\mathcal{S}$ at the cost of poorer trajectory control. Both policies are discussed in Appendix E.
@@
Algorithm 2 Hyperflux Pruning Algorithm
1:Input: Pretrained weights $\omega^{\mathrm{init}}$, pruning epochs $E_{p}$, stabilization epochs $E_{s}$ (leading to total epochs $E_{t}=E_{p}+E_{s}$), pressure scheduler Sched($s_{e},e$).
2:Output: Weights $\omega^{*}$, final topology $\mathcal{T}^{*}$.
3:Initialize:
4: Weights $\omega\leftarrow\omega^{\mathrm{init}}$ .
6: Topology $\mathcal{T}_{i}\leftarrow 1$, $\forall i\in\{1,2,...,d\}$.
7:for epoch $e=1$ to $E_{t}$ do
9:  $\omega\leftarrow\omega-\eta_{\omega}\nabla_{\omega}\mathcal{L}$.
10:  $t\leftarrow t-\eta_{t}\nabla_{t}\mathcal{J}$.
11:  if $e\leq E_{p}$ then
12:    $\gamma\leftarrow\text{{\sc Sched}}(\text{current sparsity }s_{e},e)$
13:  else
14:    $\eta_{t}\leftarrow 0.9\cdot\eta_{t}$
15:    $\gamma\leftarrow 0$
16:  end if
17:end for
4 Performance Comparison
To validate Hyperflux, we conduct comprehensive pruning experiments on a diverse set of architectures and datasets: ResNet-50 and VGG-19 on CIFAR-10/100, and ResNet-50 on ImageNet-1K. We pit Hyperflux against state-of-the-art pruning approaches such as GraNet Liu et al. (2022), GMP Zhu and Gupta (2018), Spartan Tai et al. (2022), and AC/DC Peste et al. (2021). To ensure a fair comparison, we run ourselves all other methods, initializing them with the pretrained weights used in Hyperflux, while maintaining the same training budget and augmentations. We test several training setups for each method and report the best results, to ensure no unfair degradation occurs due to suboptimal hyperparameters.
4.1 CIFAR-10 / 100

### v3

# The Neural Pruning Law Hypothesis

@@ Abstract
Network pruning is used to reduce inference latency and power consumption in large neural networks. However, most current pruning methods rely on ad-hoc heuristics that are poorly understood. We introduce Hyperflux, a conceptually-grounded pruning method, and use it to study the pruning process. Hyperflux models this process as an interaction between weight flux, the gradient’s response to the weight’s removal, and network pressure, a global regularization driving weights towards pruning. We postulate properties that arise naturally from our framework and find that the relationship between minimum flux among weights and density follows a power-law equation. Furthermore, we hypothesize the power-law relationship to hold for any effective saliency metric and call this idea the Neural Pruning Law Hypothesis. We validate our hypothesis on several families of pruning methods (magnitude, gradients, $L_{0}$), providing a potentially unifying property for neural pruning.
@@ 1 Introduction
Overparameterization has become the norm in modern deep learning to achieve state-of-the-art performance (Neyshabur et al., 2019; Allen-Zhu et al., 2019; Li et al., 2018). Despite clear benefits for training, this practice also increases computational and memory costs, complicating deployment on resource-constrained devices such as edge hardware, IoT platforms, and autonomous robots (Shi et al., 2016; Li et al., 2019). Recent theoretical and empirical findings suggest that sparse subnetworks extracted from large dense models can match or exceed the accuracy of their dense counterparts (Frankle and Carbin, 2019; Zhou et al., 2019; Ma et al., 2021; Lee et al., 2019; De Jorge et al., 2021; Cho et al., 2023; Yite et al., 2023; Frantar et al., 2024; Wang et al., 2023) and even outperform smaller dense models of equal size (Ramanujan et al., 2020; Li et al., 2020; Zhu and Gupta, 2018). These results have created interest in network pruning as a strategy to identify minimal, high-performing subnetworks.
@@
Pruning has a rich history (LeCun et al., 1989; Mozer and Smolensky, 1988; Thimm and Hoppe, 1995) and continues to prove valuable for real-time applications (Han et al., 2016; Jongsoo et al., 2017; Wang et al., 2019). Recent methods have significantly advanced the field by resorting to a variety of strategies, from magnitude pruning, gradient methods, and Hessian-based criteria (Han et al., 2015, 2016; LeCun et al., 1992; Singh and Alistarh, 2020; Bellec et al., 2018; Frankle and Carbin, 2019) to dynamic pruning approaches (Liu et al., 2020; Cho et al., 2023; Savarese et al., 2020; Kusupati et al., 2020; Wortsman et al., 2019) or combinations thereof (Liu et al., 2022; Evci et al., 2020). However, the strong interdependence between weights remains a challenge (Jin et al., 2020; Templeton et al., 2024; Lee et al., 2019; De Jorge et al., 2021; Louizos et al., 2017), as it complicates the task of determining each weight’s importance. Optimal pruning has been explored (Hoefler et al., 2021; LeCun et al., 1992), but such formulations are typically computationally intractable in practice. Overall, the field consists of a variety of seemingly unconnected pruning methods, each trying to approximate importance in its own way.
Given this gap, we ask: Are there unifying properties for pruning methods?
@@
To study these properties, we start by creating our own method, based on first principles rather than empirical results and heuristics. Inspired by the principle that the value of something is not truly known until it is lost, which has shaped major discoveries in fields such as functional genomics (Giaever et al., 2002; Shalem et al., 2014), neuroscience (Rorden and Karnath, 2004), and network science (Albert et al., 2000), we introduce Hyperflux, an $L_{0}$ pruning method that determines a weight’s importance by first removing it and observing the gradient caused as a result, its flux.
@@
The main idea of Hyperflux is that each weight has a flux, which appears when the weight is pruned through the network’s gradients. A global $L_{0}$ regularization term called pressure pushes all weights towards pruning, aiming to uncover each of their fluxes. Those weights whose flux is greater than the pressure will be regrown, while the rest will remain pruned. This process is repeated until the end of training.
We postulate and empirically confirm two properties that emerge from our framework: density convergence and a relationship betwee minimum flux among weights and density, for which we find a power-law equation in our experiments. Based on these results, we hypothesize this power-law for any saliency, including those in most pruning methods.
To confirm our hypothesis, we choose 3 established types of pruning methods, magnitude, gradient, $L_{0}$, and verify that the same power-law relationship appears for ResNet-50 and VGG-19 networks on Cifar-10, Cifar-100 and ImageNet datasets respectively. While this is not our main objective, we also observe Hyperflux to be better or competitive to current state-of-the-art methods.
@@
  • We introduce Hyperflux, a conceptually grounded pruning method that obtains competitive results compared to state-of-the-art.
  • We study Hyperflux’s mechanisms and build an analytical framework around it, leading us to a power-law relationship between minimum flux among weights and density.
  • We hypothesize the power-law relationship between minimum saliency among weights and density to hold across most pruning methods, and verify its existence on 3 families of such methods.
@@ 3.2 Flux and pressure
In Appendix A.1, we analyze Hyperflux mechanics in more detail, and show that a weight will be regrown only if the average flux is greater than the pressure for the training interval in which the weight is pruned. Since flux only exists if weights encode meaningful information, we begin pruning by initializing the network with pretrained weights.
3.3 Minimum Flux and Convergence
@@
Following from the insights about flux and pressure described so far, we postulate a series of properties that naturally emerge. We experimentally validate each of the properties, confirming our expectations, and laying the foundation for the Neural Pruning Law Hypothesis.
@@
Property 1: Density Convergence for a Fixed $\gamma$. As density decreases, fewer weights are used to represent the same information contained within the dataset. As long as the accuracy remains constant, the importance – and therefore flux – of the remaining weights should increase. Once the minimum flux among all weights, $\mathcal{W}_{f}$, matches the pressure, we expect density to converge to an equilibrium. Therefore, we ask the following question:Given a fixed $\gamma$, will the network converge to a final density $\mathcal{D}$? In Figure 2, we test this by running ResNet-50 and VGG-19 on Cifar-10, Cifar-100 respectively. We preload weights for each network and allow it to train for 1000 epochs with a constant pressure $\gamma$.
Figure 2: Convergence for fixed $\gamma=2$.
To better understand how convergence is affected by optimizers and learning, we test two different optimizers for $t$ values, SGD and Adam, along with a high and low learning rate, differing by two orders of magnitude. We observe that networks with low learning rates tend to converge faster and smoother compared to high learning rates. Further ablation studies are found in Appendix B.
Property 2: Relationship Between Minimum Weight Flux and Density. Property 1 says that all networks converge to a fixed density for a fixed $\gamma$, when $\mathcal{W}_{f}=\gamma$, and do not end up diverging or oscillating. Therefore, we ask: Is there a predictable relationship between minimum flux among weights and network density? For each experiment, we train a network with constant $\gamma$ and allow it to converge. The learning rate is kept constant, since it influences the final density (see Fig 2). Figure 3 illustrates different convergence points for several $\gamma$ values for ResNet-50, Cifar-10. Similar results can be seen in Appendix C.1 for VGG-19, Cifar-100.
Figure 3: The relationship between $\gamma$ (minimum flux) and final density for ResNet-50, Cifar-10. The 3 regions we discussed are highlighted in the figure.
We observe three regions taking shape in our experiments. In region 1, the pressure is too small for any meaningful weights to be pruned and information compression to occur. In region 2, we observe constant accuracy over a large density range, where meaningful information compression occurs, leading to a clear power-law relationship between pressure and density!
$\ln(s)\;=\;\ln(c)\;-\;\alpha_{0}\,\ln(\gamma)$ (4)
where constants, c, $\alpha_{0}$, depend on dataset and network architecture. In region 3, where accuracy starts collapsing along with the remaining weights, leading eventually to 0% density. Our Neural Pruning Law Hypothesis will be concerned with region 2.
Figure 4: The relationship between minimum saliency among weights and density, for ResNet-50 Cifar-10 on Hyperflux (representing $L_{0}$ methods), Iterative magnitude pruning and Taylor approximation (representing gradient methods). The 3 regions as well as the power-law relation are clearly delimited in all 3 cases.
3.4 The Neural Pruning Law Hypothesis
Our analysis in Section 3.2 showed that weight flux is an effective pruning criterion and that, according to Property 2, the relationship between $\mathcal{W}_{f}$ and density follows a power-law. We hypothesize that such a power-law relationship holds for ANY effective saliency measure, including most pruning methods heuristics. We denote this idea as the Neural Pruning Law Hypothesis (NPLH). While fully validating NPLH is left for future work, we provide a strong preliminary set of experiments. We include 3 types of pruning methods: magnitude pruning, represented by iterative magnitude pruning (Frankle and Carbin, 2019), gradient-based methods, represented by Taylor approximation, (Molchanov et al., 2019) and $L_{0}$ methods, represented by Hyperflux.
We run the 3 methods on combinations of Cifar-10, Cifar-100 and ImageNet datasets with ResNet-50 and VGG-19 networks, and expect to see the same 3 regions along with the power-law appear for any importance approximation. As a common baseline, we choose ResNet-50 on Cifar-10, as it is implemented in all method we chose. The training setup for NPLH is detailed in Appendix C.1. Figure 4 includes results for ResNet-50 Cifar-10 for all 3 methods. The power-law region is highlighted for each method. Further studies on VGG-19 and ResNet-50 networks trained on Cifar-100 and ImageNet are also included Appendix C.1 and display the same 3 regions.
A likely explanation for the fact that all reasonable measures of saliency follow the power law is that all these measure approximate some fundamental importance of the weights. What that importance is exactly, remains to be seen in future work.
@@ 3.5 Pressure Scheduler & Stabilization Stage
Our findings from Section 3.3 suggest that a $\gamma$ exists for any desired $\mathcal{D}$. However, in practical applications, $\gamma$ is not known at the start and tuning it would require hyperparameter search. Instead, we propose a dynamic scheduler that adjusts $\gamma$ after each epoch automatically, driving the network towards a desired sparsity. Furthermore, to ensure convergence of weights after pruning, we introduce a stabilization stage at the end.
@@
1:Input: Current sparsity $d_{e}$ and epoch $e$
2:Requires: Pruning epochs $E_{p}$, desired final sparsity $\mathcal{D}$, pressure policy $\Pi$, step $u$, exponent $\alpha$.
@@
Pressure Scheduler: The goal of our scheduler is to adjust $\gamma$ such that the network converges to $\mathcal{D}$ with minimal accuracy decrease. We denote by $\gamma_{e}$ and $d_{e}$, the pressure and network density at epoch $e$. Because the frequency of updates is constant, occurring after each epoch, any non-linear change in pressure required to affect density (see Eq. 4) must arise from the update rule. To get this nonlinearity we set $\gamma_{e}=(p_{e})^{\alpha},$ with $p_{e}$, a scalar base, updated according to Algorithm 1. Inertia terms $p_{+}$ and $p_{-}$ account for suboptimal $\alpha$ or $u$. Apart from non-linear updates, our scheduler requires a binary pressure policy $\Pi$ to determine when those updates are applied such that $\mathcal{D}$ is reached. We explore two choices for $\Pi$. In the first, $d_{e}$ follows a user-defined curve $f(e)$, trading precise convergence to $\mathcal{D}$ for trajectory control. In this case $\Pi(E_{p},\mathcal{D},d_{e},e)$ is true if and only if $d_{e}<f(e)$. In the second policy, $d_{e}$ stays between a dynamic upper bound and the target $\mathcal{D}$, achieving precise convergence to $\mathcal{D}$ at the cost of poorer trajectory control. Both policies are discussed in Appendix E.
@@
Algorithm 2 Hyperflux Pruning Algorithm
1:Input: Pretrained weights $\omega^{\mathrm{init}}$, pruning epochs $E_{p}$, stabilization epochs $E_{s}$ (leading to total epochs $E_{t}=E_{p}+E_{s}$), pressure scheduler Sched($d_{e},e$).
2:Output: Weights $\omega^{*}$, final topology $\mathcal{T}^{*}$.
3:Initialize:
4: Weights $\omega\leftarrow\omega^{\mathrm{init}}$ .
6: Topology $\mathcal{T}_{i}\leftarrow 1$, $\forall i\in\{1,2,...,d\}$.
7:for epoch $e=1$ to $E_{t}$ do
9:  $\omega\leftarrow\omega-\eta_{\omega}\nabla_{\omega}\mathcal{L}$.
10:  $t\leftarrow t-\eta_{t}\nabla_{t}\mathcal{J}$.
11:  if $e\leq E_{p}$ then
12:   $\gamma\leftarrow\text{{\sc Sched}}(\text{current density }d_{e},e)$
13:  else
14:   $\eta_{t}\leftarrow 0.9\cdot\eta_{t}$
15:   $\gamma\leftarrow 0$
16:  end if
17:end for
4 Hyperflux Performance
We now show that Hyperflux, in addition to allowing us to come up with NPLH, also has strong empirical performance. To validate Hyperflux, we conduct comprehensive pruning experiments on a diverse set of architectures and datasets: ResNet‑50 and VGG‑19 on CIFAR‑10/100, and ResNet‑50 on ImageNet‑1K. We pit Hyperflux against state-of-the-art pruning approaches such as GraNet Liu et al. (2022), GMP Zhu and Gupta (2018), Spartan Tai et al. (2022), and AC/DC Peste et al. (2021). To ensure a fair comparison, we run ourselves all other methods, initializing them with the pretrained weights used in Hyperflux, while maintaining the same training budget and augmentations. We test several training setups for each method and report the best results, to ensure no unfair degradation occurs due to suboptimal hyperparameters.
4.1 CIFAR-10 / 100